# Observation Interference in Partially Observable Assistance Games

### v1

@@ 4.1 Revealing errors can emerge as an optimal POAG solution
Past work has shown how RLHF can cause misleading [35] and deceptive [37, 22] behaviors. Specifically, Lang et al. [22] show that in order to get better human feedback, RLHF can have an incentive to hide error messages. In contrast, we show with the following example that revealing error messages can emerge in POAG solutions.
@@
Why does RLHF have an incentive to hide error messages, while the POAG solution has an incentive to reveal the errors? In RLHF, the agent is merely maximizing the feedback it receives from the human, rather than the human’s true reward function. If an RLHF agent can deceive the human to get better feedback, it has an incentive to do so. In contrast, optimal POAG agents only care about the human’s true reward and will reveal errors when that information is useful to the human.
@@ 4.2 Communicating private information is an incentive for observation interference at the action level
Intuitively, one might hope that $\mathbf{A}$ would never take observation-interfering actions. After all, classic theory tells us that when $\mathbf{H}$ is in a single-agent setting, the value of perfect information is nonnegative: more informative observations never lead to worse solutions. But as it turns out, when $\mathbf{H}$ and $\mathbf{A}$ interact in a POAG, there are cases in which all optimal policy pairs require $\mathbf{A}$ to take observation-interfering actions. The main reason for $\mathbf{A}$ to take observation-interfering actions is to communicate its own private information to $\mathbf{H}$. Consider the following example.
@@
First, $\mathbf{A}$ takes an action. For each of the 10 total cuda versions, $\mathbf{A}$ can choose to or not to suppress it from the list of available packages. This gives $\mathbf{A}$ $2^{10}$ total actions, where 1 action is non-observation interference (suppressing nothing), and the remaining $12^{10}-$ 1 actions interfere with observations.
@@
The need for $\mathbf{A}$ to take observation-interfering actions to communicate to $\mathbf{H}$ disappears if $\mathbf{A}$ has other means of communication. For instance, if in Example 4.3, $\mathbf{A}$ could simply tell $\mathbf{H}$ what to do, then $\mathbf{A}$ wouldn’t need to interfere. To formalize this intuition, we now prove that if $\mathbf{A}$ can communicate with $\mathbf{A}$, then there is always an optimal policy pair that does not require interference.
@@
Let $M$ be any POAG, and provide $\mathbf{A}$ with an unbounded communication channel to $\mathbf{H}$, forming $M^{\mathbf{A}\rightarrow\mathbf{H}}$. Then there is an optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$ for $M^{\mathbf{A}\rightarrow\mathbf{H}}$ where $\pi^{\mathbf{A}}$ does not interfere with observations at the action level (and $\pi^{\mathbf{H}}$ observes naively).
Note that under the conditions in the theorems $\mathbf{H}$ may still need to act non-naively in order to communicate her private information to $\mathbf{A}$ (as shown in Section 5) [cf. 1].
@@ 4.3 Optimal policy pairs never require observation interference at the policy level
Compared to our previous action-level notion of observation interference (Definition 3.2), this new policy-level notion (Definition 4.6) differs in how it treats $\mathbf{H}$’s inference process. Whereas the action-level notion models inference about isolated observations, the policy-level notion allows $\mathbf{H}$ to make inferences in the context of $\mathbf{A}$’s overall strategy. In this broader framework, cases which appear to destroy information when viewed at the action level may actually provide new information when viewed at the policy level. In fact, we show in the following theorem that it’s never strictly necessary to interfere with observations at the policy level.
@@
We now revisit Example 4.3. When just considering $\mathbf{H}$’s observations in isolation, seeing the list of all available cuda versions is strictly more informative than having some of the available versions suppressed. Suppose, however, that $\mathbf{H}$ knows $\mathbf{A}$’s policy is to filter the list by suppressing only the incompatible versions. Then, compared to seeing the list of all available versions, receiving the filtered list provides new information. $\mathbf{H}$’s ability to infer information based on knowledge of $\mathbf{A}$’s policy is what motivates Definition 4.6. Accordingly, when $\pi^{\mathbf{A}}$ is filtering the list by suppressing only the non-compatible versions, $\mathbf{A}$ is interfering with observations at the action level but not at the policy level.
@@ 5 Querying $\mathbf{H}$’s preferences is an incentive for observation interference
At first sight, this may appear to be a counterexample to Theorem 4.2. However, note that Example 5.1 actually does have optimal policy pairs in which $\mathbf{A}$ doesn’t interfere. In particular, even if $\mathbf{A}$ does not interfere and the two available nodes are CPU-optimized, $\mathbf{H}$ may simply communicate her CPU-versus-GPU preference anyway! That is, when facing a choice between CPU-optimized node 1 and 2, she may choose, say, 1 if she favors GPU-optimized nodes and 2 if she favors CPU-optimized nodes. However, this type of human strategy seems implausible, as it would require $\mathbf{H}$ and $\mathbf{A}$ to have settled on some communication strategy that overrides $\mathbf{H}$’s immediate preferences about the machines that $\mathbf{H}$ can in fact choose between.
@@
Using our notion of acting naively (Definition 3.7), we state the following result (with proof in Appendix C):
@@
There is a POAG $M$ with the following properties. For every optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$, at least one of these holds:
  1. i)
    $\pi^{\mathbf{H}}$ is not acting naively, or
  2. ii)
    $\pi^{\mathbf{A}}$ interferes with observations at both the action and policy levels.
Additionally, there exists an optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$ where $\pi^{\mathbf{H}}$ acts naively and $\pi^{\mathbf{A}}$ interferes with observations at both the action and policy levels.
@@
Intuitively, the problem in the above example is that the human has private information that she needs to communicate with her choices. (Because her choices yield different immediate rewards, naive choices fail to communicate.) As before, the need for interference or non-naive choice disappears if the human has no private information to provide. Since in a POAG, we assume that $\mathbf{H}$ always has at least some private information about her preferences $\theta$, we omit a formal result. The following shows that the need for interference / non-naivete also disappears if $\mathbf{H}$ can communicate with $\mathbf{A}$. To also rule out the need to interfere with observations for $\mathbf{A}\rightarrow\mathbf{H}$ communication (discussed in Section 4.2) we assume communication channels in both direction.
@@ 6 Human irrationality is an incentive for observation interference
Finally we consider a third reason for observation interference: human irrationality or bounded rationality. Roughly, reducing the amount of information supplied to the human may simplify the human’s decision problem and thus improve her decision making. Importantly, this motivation for observation interference may exist even if neither $\mathbf{H}$ nor $\mathbf{A}$ has any private information.
As our model of human decision making, we adopt Boltzmann rationality [24, 25], which has recently been used in (C)IRL [21, 29, 39]. We define Boltzmann rationality as follows:
@@
Mathematically speaking, a Boltzmann-rational agent at each time step computes the expected utilities of each of the available actions and then randomizes according to the softmax of the expected utilities.
The central feature of the Boltzmann rationality model is that it postulates that agents are more likely to get decisions right if the differences in expected utility of the options are large. It’s easy to see that if the human observes naively (and thus doesn’t have calibrated beliefs), $\mathbf{A}$ sometimes prefers observation interference. Roughly, $\mathbf{A}$ wants to make $\mathbf{H}$ always believe that the difference in utilities between her actions is high.
However, it turns out that even if the Boltzmann-rational human has calibrated beliefs, $\mathbf{A}$’s optimal policy sometimes interferes with observations, even if neither $\mathbf{A}$ nor $\mathbf{H}$ has private information. Intuitively, providing more information may sometimes result in less clear-cut decisions, i.e., decision situations with a smaller difference between the correct and incorrect option. To illustrate this phenomenon, consider the following example.
@@
$\mathbf{H}$ is running a terminal command and is unsure whether to run the command with flag 1 or flag 2. With equal probability, either flag 1 or flag 2 is better, and how good the flags are differs by either a little or a lot. Thus, $\mathbf{H}$ is uniformly at random in one of four states. $\mathbf{A}$ has two actions: man and tldr. The man page is a long document that tells the human exactly what the values of the flags are (ie, exactly what state the human is in). The tldr page is a short summary that tells the human which flag is better, but not by how much (ie, ruling out half the states, leaving half remaining).
Intuitively, both the tldr and man pages allow the human to choose optimally, but the man page is more complicated and therefore more likely to be misinterpreted. Choosing specific utilities, the effect of interference under Boltzmann rationality is as follows. If $\mathbf{A}$ interferes (i.e., provides the tldr page), then $\mathbf{H}$ always chooses between a utility of $4$ and 00. If $\mathbf{A}$ does not interfere, then half the time, $\mathbf{H}$ chooses between utilities and 1 and 0, and half the time $\mathbf{H}$ chooses between utilities 7 and 0. It turns out that for $\beta=1$, $\mathbf{H}$ achieves higher utility in expectation under the condition where $\mathbf{A}$ interferes. Building on this idea, we can prove the following (with details in Appendix D).
@@
One might expect that the need for interference is greater at smaller values of $\beta$ and disappears at larger values of $\beta$. After all, we know by Theorem 4.2 that $\mathbf{A}$ has no need for observation interference when $\mathbf{H}$ is perfectly rational. However, it turns out that this is not the case! For instance, in Example 6.2 (with the above numbers), $\mathbf{A}$ prefers interference if (and only if) $\beta$ is above $\approx 0.77361$. Roughly speaking, the reason is that at low values of $\beta$, the observation’s effect is dominated by getting the $s_{a}$/$s_{c}$ case right more often. At high values of $\beta$, the observation’s effect is dominated by getting the $s_{b}$/$s_{d}$ case right less often.
@@ 7 Experiments
In the previous sections, we explored why AI assistants might take observation-interfering actions. Section 4 showed that sometimes they interfere with observations at the action level in order to communicate other, more important information at the policy level. Section 6 showed that sometimes they interfere with observations to make decisions easier for humans. Now, we develop a model game to analyze these behaviors. We run experiments to answer the following questions within our model:
@@ 7.1 Experiment details
We consider a game with $d=5$ products. We vary $R$’s number of interferences $k\in\{0,1,2,3,4\}$. We run a Monte Carlo simulation with $30{,}000$ trials to calculate the expected payoff in each setting. We run our experiments with a CPU runtime on Google Colab.
@@ 7.2 Varying $\mathbf{H}$’s rationality
(a) $\mathbf{A}$’s Number of Private Observations = 2
(b) $\mathbf{H}$’s Rationality Coefficient $\beta=\infty$
@@
When $\mathbf{H}$ is highly irrational at $\beta=0.01$, $\mathbf{A}$ should interfere with as many sensors as possible. This effectively lets $\mathbf{A}$ choose $\mathbf{H}$’s action. When $\mathbf{H}$ is acting little better than randomly, it’s best for $\mathbf{A}$ to choose $\mathbf{H}$’s action, even when $\mathbf{A}$ has less information than $\mathbf{H}$. For larger values of $\beta$, a tradeoff emerges. As $\mathbf{A}$ has two private observations, there is an increasing benefit to interfere to communicate information to $\mathbf{H}$. However, as $\mathbf{H}$ can now make use of their own private observations, $\mathbf{A}$ must be careful not to destroy too much of $\mathbf{H}$’s private information by excessive interference.

### v2

@@ 4.1 Revealing Errors can Emerge as an Optimal POAG Solution
We can model RLHF within the POAG framework as follows: $\mathbf{A}$’s goal in RLHF is to satisfy $\mathbf{H}$’s preferences. In a POAG, this corresponds to the shared reward function $R$ which has a parameterization $\theta$ that only $\mathbf{H}$ knows. In RLHF, $\mathbf{A}$ rolls out trajectories, and $\mathbf{H}$ picks which trajectory is preferred. A POAG can model this by letting $\mathbf{H}$ observe pairs of trajectories explored by $\mathbf{A}$ but only giving $\mathbf{H}$ a binary action (to choose which trajectory $\mathbf{H}$ prefers). In RLHF, $\mathbf{A}$’s final policy maximizes an estimate of $R$ based on a dataset of $\mathbf{H}$’s preference comparisons (Lang et al., 2024, Proposition 4.1). In the POAG framework, $\mathbf{A}$ can compute this policy based on $\mathbf{A}$’s observations of $\mathbf{H}$’s binary actions.
Past work has shown how RLHF can cause misleading (Wen et al., 2024) and deceptive (Williams et al., 2024; Lang et al., 2024) behaviors. Specifically, Lang et al. (2024, Example B.1) show that in order to get better human feedback, RLHF can have an incentive to hide error messages.
In contrast to RLHF, we show with the following example that revealing error messages can emerge in POAG solutions.
@@ 4.2 Communicating Private Information is an Incentive for Observation Interference at the Action Level
One might hope that $\mathbf{A}$ would never take observation-interfering actions. After all, classic theory tells us that when $\mathbf{H}$ is in a single-agent setting, the value of (perfect) information is nonnegative (e.g., Koller & Friedman, 2009, Sect. 23.7; Russell & Norvig, 2010, Sect. 16.6.3): more informative observations never lead to worse solutions. But as it turns out, when $\mathbf{H}$ and $\mathbf{A}$ interact, there are some POAGs in which all optimal policy pairs require $\mathbf{A}$ to take observation-interfering actions. The main reason for $\mathbf{A}$ to take observation-interfering actions is to communicate its own private information to $\mathbf{H}$. Consider the following example.
@@
First, $\mathbf{A}$ takes an action. For each of the 10 total cuda versions, $\mathbf{A}$ can choose to or not to suppress it from the list of available packages. This gives $\mathbf{A}$ $2^{10}$ total actions, where 1 action is non-observation interference (suppressing nothing), and the remaining $2^{10}-1$ actions interfere with observations.
@@
The need for $\mathbf{A}$ to take observation-interfering actions to communicate to $\mathbf{H}$ disappears if $\mathbf{A}$ has other means of communication. For instance, if in Example 4.3, $\mathbf{A}$ could simply tell $\mathbf{H}$ what to do, then $\mathbf{A}$ wouldn’t need to interfere. To formalize this intuition, we now prove that if $\mathbf{A}$ can communicate with $\mathbf{H}$, then there is always an optimal policy pair that does not require interference.
@@
Let $M$ be any POAG, and provide $\mathbf{A}$ with an unbounded communication channel to $\mathbf{H}$, forming $M^{\mathbf{A}\rightarrow\mathbf{H}}$. Then there is an optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$ for $M^{\mathbf{A}\rightarrow\mathbf{H}}$ where $\pi^{\mathbf{A}}$ does not interfere with observations at the action level and $\pi^{\mathbf{H}}$ observes naively.
One could argue that in practice, an unrestricted communication channel between $\mathbf{A}$ and $\mathbf{H}$ could usually be made available. However, Theorem 4.5 ignores various real-world obstacles. For one, it considers communication that incurs no cost, but in reality, communication costs $\mathbf{H}$ time and effort. Second, the optimal policy pair requires $\mathbf{A}$ to send information in a way that $\mathbf{H}$ can reliably understand and act upon. We expect that in practice, $\mathbf{A}$ and $\mathbf{H}$ sometimes cannot understand each other. Therefore, despite Theorem 4.5, we think observation interference is of broad practical relevance, even where $\mathbf{A}$ can, e.g., send text messages to $\mathbf{H}$.
@@ 4.3 Optimal Policy Pairs Never Require Observation Interference at the Policy Level
Compared to our previous action-level notion of observation interference (Definition 3.2), this new policy-level notion (Definition 4.6) differs in how it treats $\mathbf{H}$’s inference process. Whereas the action-level notion models inference about isolated observations, the policy-level notion allows $\mathbf{H}$ to make inferences in the context of $\mathbf{A}$’s overall strategy.
We now revisit Example 4.3. For observation tampering under our earlier Definition 3.2, $\mathbf{H}$ simply knows that $\mathbf{A}$ has taken the action to suppress some versions of cuda; $\mathbf{H}$ does not know anything about $\mathbf{A}$’s policy. For all $\mathbf{H}$ knows, $\mathbf{A}$’s policy could be to randomly suppress cuda versions or to always suppress the same cuda version. Thus, suppressing any version is strictly less informative for $\mathbf{H}$ than the list of all available versions. This is why Definition 3.2 calls suppressing versions “tampering at the action level.”
The key difference with Definition 4.6 is that $\mathbf{H}$ knows $\mathbf{A}$’s policy. Suppose that $\mathbf{A}$’s policy $\pi^{\mathbf{A}}$ is to suppress exactly the versions of cuda that are incompatible with the other software in the environment. Because $\mathbf{H}$ knows that $\mathbf{A}$ suppressed the incompatible cuda versions, seeing the filtered list tells $\mathbf{H}$ which versions of cuda are compatible! Although suppressing versions is strictly less informative under Definition 3.2 (when $\mathbf{H}$ doesn’t know $\mathbf{A}$’s policy), suppressing versions provides $\mathbf{H}$ with new information under Definition 4.6 (when $\mathbf{H}$ knows $\mathbf{A}$’s policy). Accordingly, $\pi^{\mathbf{A}}$ is interfering with observations at the action level but not at the policy level.
In our examples, we will mostly consider actions that in some sense act directly on $\mathbf{H}$’s observations. Yet Definition 4.6 also considers the informational effects of physical actions. For example, if $\mathbf{A}$ (visibly) tries to open a door that $\mathbf{A}$ knows to be locked, then this reveals to $\mathbf{H}$ that the door is locked. Consequently, not trying to open the door (when $\mathbf{A}$ knows it to be locked) is an instance of observation interference in the sense of Definition 4.6. While having the same (null) effect on the state of the world, trying to open the door provides $\mathbf{H}$ with more information.
As in Example 4.3, cases which appear to destroy information when viewed at the action level may actually provide new information when viewed at the policy level. In fact, the following theorem shows that it’s never strictly necessary to interfere with observations at the policy level.
@@ 5 Querying $\mathbf{H}$’s Preferences is an Incentive for Observation Interference
At first sight, Example 5.1 may appear to be a counterexample to Theorem 4.2, which states that a non-interfering optimal policy pairs exists. However, note that Example 5.1 actually does have optimal policy pairs in which $\mathbf{A}$ doesn’t interfere. In particular, even if $\mathbf{A}$ does not interfere and the two available nodes are CPU-optimized, $\mathbf{H}$ may simply communicate her CPU-versus-GPU preference anyway! That is, when facing a choice between CPU-optimized node 1 and 2, she may choose, say, 1 if she favors GPU-optimized nodes and 2 if she favors CPU-optimized nodes. However, this type of human strategy seems implausible, as it would require $\mathbf{H}$ and $\mathbf{A}$ to have settled on some communication strategy that overrides $\mathbf{H}$’s immediate preferences about the machines that $\mathbf{H}$ can in fact choose between.
The key point of Example 5.1 is that—while there is some optimal policy pair without observation interference—there is no plausible optimal policy pair that avoids observation interference. More specifically, we use the notion of acting naively (Definition 3.7) to express this notion of plausibility and rule out the above policy. We thus obtain the following proposition (with proof in Appendix C): in some POAGs, if we want to play an optimal policy pair and we want $\mathbf{H}$ to be able to act naively, then $\mathbf{A}$ has to interfere with observations.
@@
There is a POAG $M$ with the following properties. For every optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$, at least one of the following holds: (i) $\pi^{\mathbf{H}}$ is not acting naively, or (ii) $\pi^{\mathbf{A}}$ interferes with observations at both the action and policy levels. Additionally, there exists an optimal policy pair $(\pi^{\mathbf{H}},\pi^{\mathbf{A}})$ where $\pi^{\mathbf{H}}$ acts naively and $\pi^{\mathbf{A}}$ interferes with observations at both the action and policy levels.
@@
Intuitively, the problem in the above example is that the human has private information that she needs to communicate with her choices. (Because her choices yield different immediate rewards, naive choices fail to communicate.) As before, the need for interference or non-naive choice disappears if the human has no private information to provide. The following shows that the need for interference / non-naivete also disappears if $\mathbf{H}$ can communicate with $\mathbf{A}$. To also rule out the need to interfere with observations for $\mathbf{A}\rightarrow\mathbf{H}$ communication (discussed in Section 4.2) we assume communication channels in both direction.
@@ 6 Human Irrationality is an Incentive for Observation Interference
Finally, we consider a third reason for observation interference: human irrationality or bounded rationality. Roughly, reducing the amount of information supplied to the human may simplify the human’s decision problem and thus improve her decision making. Importantly, this motivation for observation interference may exist even if neither $\mathbf{H}$ nor $\mathbf{A}$ has any private information.
As our model of human decision making, we adopt Boltzmann rationality (Luce, 1959; McFadden, 1973), which has recently been used in (C)IRL (Laidlaw & Dragan, 2021; Ramachandran & Amir, 2007; Ziebart et al., 2008).
@@
It turns out that even if the Boltzmann-rational human has calibrated beliefs, $\mathbf{A}$’s optimal policy sometimes interferes with observations, even if neither $\mathbf{A}$ nor $\mathbf{H}$ has private information. Intuitively, providing more information may sometimes result in less clear-cut decisions, i.e., decision situations with a smaller difference between the correct and incorrect option. To illustrate this phenomenon, consider the following example.
@@
$\mathbf{H}$ is running a terminal command and is unsure whether to run the command with flag 1 or flag 2. With equal probability, either flag 1 or flag 2 is better, and how good the flags are differs by either a little or a lot. The worse flag always yields a utility of 0, while the better flag either yields a utility of 1 or a utility 7. Thus, $\mathbf{H}$ is uniformly at random in one of four states. $\mathbf{A}$ has two actions: man and tldr. The man page is a long document that tells the human exactly what the values of the flags are (i.e., the exact state: which flag is better and whether its utility is 1 or 7). The tldr page is a short summary that tells the human which flag is better, but not by how much (i.e., ruling out half the states, leaving half remaining). Thus, the expected utility of the better flag is 4 (and of the worse flag is 0).
Intuitively, both the tldr and man pages allow the human to choose optimally, but the man page is more complicated and therefore more likely to be misinterpreted. Choosing specific utilities, the effect of interference under Boltzmann rationality is as follows. If $\mathbf{A}$ interferes (i.e., provides the tldr page), then $\mathbf{H}$ always chooses between a utility of $4$ and 0. If $\mathbf{A}$ does not interfere, then half the time, $\mathbf{H}$ chooses between utilities 1 and 0, and half the time $\mathbf{H}$ chooses between utilities 7 and 0. It turns out that for $\beta=1$, $\mathbf{H}$ achieves higher utility in expectation under the condition where $\mathbf{A}$ interferes. Building on this idea, we can prove the following (with details in Appendix D).
@@ 7 Experiments
Motivated by the theory in Section 4 and Section 6, we develop a model game and run experiments to analyze:
@@ 7.1 Experiment Details
(a) $\mathbf{A}$’s Number of Private Observations = 2
(b) $\mathbf{H}$’s Rationality Coefficient $\beta=\infty$
@@
We consider a game with $d=5$ products. We vary $R$’s number of interferences $k\in\{0,1,2,3,4\}$. We run a Monte Carlo simulation with 30,00030{,}000 trials to calculate the expected payoff in each setting.
@@ 7.2 Varying $\mathbf{H}$’s Rationality
When $\mathbf{H}$ is highly irrational at $\beta=0.01$, $\mathbf{A}$ should interfere with as many sensors as possible, effectively choosing $\mathbf{H}$’s action. When $\mathbf{H}$ is acting little better than randomly, it’s best for $\mathbf{A}$ to choose $\mathbf{H}$’s action, even when $\mathbf{A}$ has less information than $\mathbf{H}$. For larger values of $\beta$, a tradeoff emerges. As $\mathbf{A}$ has two private observations, there is an increasing benefit to interfere to communicate information to $\mathbf{H}$. However, as $\mathbf{H}$ can now make use of their own private observations, $\mathbf{A}$ must be careful not to destroy too much of $\mathbf{H}$’s private information by excessive interference.